OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) and has a slope of 3 in point-slope form. x + 1 = 3(y – 5) x – 5 = 3(y + 1) y – 5 = 3(x + 1) y + 1 = 3(x – 5)

Vocaloid (vocaloid):

point slope form: y - y1 = m(x-x1) where your point (-1,5) = (x1,y1) and m = slope